# Prioritized Information Bottleneck Theoretic Framework with Distributed Online Learning for Edge Video Analytics

### v1

@@ VI Conclusion
In this paper, we have proposed the Prioritized Information Bottleneck (PIB) framework as a robust solution for collaborative edge video analytics. Our contributions are two-fold. First, we have developed a prioritized inference mechanism to intelligently determine the importance of different camera’ FOVs, effectively addressing the constraints imposed by channel capacity and data redundancy. Second, the PIB framework showcases its effectiveness by notably decreasing communication overhead and improving tracking accuracy without requiring video reconstruction at the edge server. Extensive numerical results show that: PIB not only surpasses the performance of conventional methods like TOCOM-TEM, JPEG, H.264, H.265, and AV1 with a marked improvement of up to 17.8% in MODA but also achieves a considerable reduction in communication costs by 82.80%, while retaining low latency and high-quality multi-view sensory data processing under less favorable channel conditions.

### v2

@@ VI Conclusion
In this paper, we have proposed the Prioritized Information Bottleneck (PIB) framework as a robust solution for collaborative edge video analytics. Our contributions are two-fold. First, we have developed a prioritized inference mechanism to intelligently determine the importance of different camera’ FOVs, effectively addressing the constraints imposed by channel capacity and data redundancy. Second, the PIB framework showcases its effectiveness by notably decreasing communication overhead and improving tracking accuracy without requiring video reconstruction at the edge server. Extensive real-world experiments show that: PIB not only surpasses the performance of conventional methods like TOCOM-TEM, JPEG, H.264, H.265, and AV1 with a marked improvement of up to 17.8% in MODA but also achieves a considerable reduction in communication costs by 82.65%, while retaining low latency and high-quality multi-view sensory data processing under less favorable channel conditions.
Appendix A The Network Architecture of PIB
The PIB framework is designed with efficient computational distribution across the camera and edge server to achieve low latency and high accuracy. As illustrated in Figs. 4 and 4, the detailed network architecture is given as follows:
Camera: The camera executes the first two stages of the pipeline: (i) Feature Extraction and (ii) Hyper Encoder. These stages preprocess the raw video data into a compressed intermediate representation suitable for transmission to the edge server.
Edge Server: Upon receiving the compressed bitstream, the edge server executes (iii) Hyper Decoder, (iv) Projection and Multiview Aggregation, and (v) Spatial Aggregation and Classification. These stages reconstruct the feature maps, fuse multiview information, and generate the pedestrian occupancy map. Below is the detailed breakdown of each stage:
(i) Feature Extraction (ResNet-18 Backbone): The feature extraction employs a modified ResNet-18 backbone to retain spatial resolution critical for subsequent projection and fusion. We assume that $B$ denotes the batch size, $H$ and $W$ are the height and width of the input image.
$\begin{array}[]{|c|c|c|}\hline\cr\text{Layer Name}&\text{Input Dimensions}&\text{Output Dimensions}\\ \hline\cr\text{Input Image}&[B,3,H,W]&[B,3,720,1280]\\ \text{ResNet-18 (Part 1)}&[B,3,720,1280]&[B,64,180,320]\\ \text{ResNet-18 (Part 2)}&[B,64,180,320]&[B,512,90,160]\\ \text{Feature Extraction}&[B,512,90,160]&[B,8,90,160]\\ \hline\cr\end{array}$
(ii) Hyper Encoder and (iii) Decoder for Compression: The Hyper Encoder compresses the extracted features at the camera, while the Hyper Decoder reconstructs them at the edge server.
$\begin{array}[]{|c|c|c|}\hline\cr\text{Layer Name}&\text{Input Dimensions}&\text{Output Dimensions}\\ \hline\cr\text{Hyper Encoder}&[B,8,90,160]&[B,4,30,40]\\ \text{Hyper Decoder}&[B,4,30,40]&[B,8,90,160]\\ \hline\cr\end{array}$
(iv) Projection and Multiview Aggregation: Feature maps are projected onto a common ground plane and aggregated with coordinate maps for multiview fusion. $H_{g}$ and $W_{g}$ are the height and width of the projected ground plane grid.
$\begin{array}[]{|c|c|c|}\hline\cr\text{Layer Name}&\text{Input Dimensions}&\text{Output Dimensions}\\ \hline\cr\text{Projection}&[B,8,90,160]&[B,8,H_{g},W_{g}]\\ \text{Concatenation}&[B,8,H_{g},W_{g}]&[B,N\times 8+2,H_{g},W_{g}]\\ \hline\cr\end{array}$
(v) Spatial Aggregation and Classification: The aggregated features are processed to produce the final pedestrian occupancy map.
$\begin{array}[]{|c|c|c|}\hline\cr\text{Layer Name}&\text{Input Dimensions}&\text{Output Dimensions}\\ \hline\cr\text{Map Classifier}&[B,N\times 8+2,H_{g},W_{g}]&[B,1,H_{g},W_{g}]\\ \hline\cr\end{array}$